### v3

# Cop-Width, Flip-Width and Strong Colouring Numbers

Abstract
Cop-width and flip-width are new families of graph parameters introduced by Toruńczyk (2023) that generalise treewidth, degeneracy, generalised colouring numbers, clique-width and twin-width. In this paper, we bound the cop-width and flip-width of a graph by its strong colouring numbers. In particular, we show that for every $r\in\mathbb{N}$, every graph $G$ has $\operatorname{copwidth}_{r}(G)\leq\operatorname{scol}_{4r}(G)$. This implies that every class of graphs with linear strong colouring numbers has linear cop-width and linear flip-width. We use this result to deduce improved bounds for cop-width and flip-width for various sparse graph classes.
keywords:
cop-width, flip-width, colouring number, bounded expansion, sparsity
1 Introduction
Kierstead and Yang (2003) introduced the following definitions [^1] We consider simple, finite, undirected graphs $G$ with vertex-set ${V(G)}$ and edge-set ${E(G)}$. See (Diestel, 2017) for graph-theoretic definitions not given here. Let :=${\mathbb{N}:=\{1,2,\dots\}}$. For integers $a,b$ where $a\leq b$, let $[a,b]:=\{a,a+1,\dots,b-1,b\}$ and for $n\in\mathbb{N}$, let $[n]:=[1,n]$. A graph class is a collection of graphs closed under isomorphism. For a graph $G$ and a vertex $v\in V(G)$, let ${N_{G}(v)}:=\{w\in V(G)\colon vw\in E(G)\}$.. For a graph $G$, a total order $\preceq$ of $V(G)$, a vertex $v\in V(G)$, and an integer $r\geq 1$, let ${R(G,\preceq,v,r)}$ be the set of vertices $w\in V(G)$ for which there is a path $(v=w_{0},w_{1},\dots,w_{r^{\prime}}=w)$ of length $r^{\prime}\in[0,r]$ such that $w\preceq v$ and $v\prec w_{i}$ for all $i\in[r^{\prime}-1]$, and let ${Q(G,\preceq,v,r)}$ be the set of vertices $w\in V(G)$ for which there is a path $v=w_{0},w_{1},\dots,w_{r^{\prime}}=w$ of length $r^{\prime}\in[0,r]$ such that $w\preceq v$ and $w\prec w_{i}$ for all $i\in[r^{\prime}-1]$. For a graph $G$ and integer $r\geq 1$, the $r$-strong colouring number of $G$, $\operatorname{scol}_{r}(G)$, is the minimum integer such that there is a total order $\preceq$ of $V(G)$ with $|R(G,\preceq,v,r)|\leq\operatorname{scol}_{r}(G)$ for every vertex $v$ of $G$. Likewise, the $r$-weak colouring number of $G$, $\operatorname{wcol}_{r}(G)$, is the minimum integer such that there is a total order $\preceq$ of $V(G)$ with $|Q(G,\preceq,v,r)|\leq\operatorname{wcol}_{r}(G)$ for every vertex $v$ of $G$.
Generalised colouring numbers provide upper bounds on several graph parameters of interest. First note that $\operatorname{scol}_{1}(G)=\operatorname{wcol}_{1}(G)$ which equals the degeneracy of $G$ plus 1, implying $\chi(G)\leq\operatorname{scol}_{1}(G)$. A proper graph colouring is acyclic if the union of any two colour classes induces a forest; that is, every cycle is assigned at least three colours. For a graph $G$, the acyclic chromatic number of $G$, $\chi_{\text{a}}(G)$, is the minimum integer $k$ such that $G$ has an acyclic $k$-colouring. Kierstead and Yang (2003) proved that $\chi_{\text{a}}(G)\leq\operatorname{scol}_{2}(G)$ for every graph $G$. Other parameters that can be bounded by strong and weak colouring numbers include game chromatic number (Kierstead and Trotter, 1994; Kierstead and Yang, 2003), Ramsey numbers (Chen and Schelp, 1993), oriented chromatic number (Kostochka et al., 1997), arrangeability (Chen and Schelp, 1993), boxicity (Esperet and Wiechert, 2018), odd chromatic number (Hickingbotham, 2023) and conflict-free chromatic number (Hickingbotham, 2023).
Another attractive aspect of strong colouring numbers is that they interpolate between degeneracy and treewidth222 A tree-decomposition of a graph $G$ is a collection $\mathcal{W}=(B_{x}\colon x\in V(T))$ of subsets of $V(G)$ indexed by the nodes of a tree $T$ such that (i) for every edge $vw\in E(G)$, there exists a node $x\in V(T)$ with $v,w\in B_{x}$; and (ii) for every vertex $v\in V(G)$, the set $\{x\in V(T)\colon v\in B_{x}\}$ induces a (connected) subtree of $T$. The width of $\mathcal{W}$ is $\max\{\lvert B_{x}\rvert\colon x\in V(T)\}-1$. The treewidth $\operatorname{tw}(G)$ of a graph $G$ is the minimum width of a tree-decomposition of $G$. . As previously noted, $\operatorname{scol}_{1}(G)$ equals the degeneracy of $G$ plus $1$. At the other extreme, Grohe et al. (2018) showed that $\operatorname{scol}_{r}(G)\leq\operatorname{tw}(G)+1$ for every $r\in\mathbb{N}$, and indeed $\operatorname{scol}_{r}(G)\to\operatorname{tw}(G)+1$ as $r\to\infty$.
Generalised colouring numbers are important because they characterise bounded expansion (Zhu, 2009) and nowhere dense classes (Grohe et al., 2018), and have several algorithmic applications Dvořák (2013); Grohe et al. (2017). Let $G$ be a graph and $r\geq 0$ be an integer. A graph $H$ is an $r$-shallow minor of $G$ if $H$ can be obtained from a subgraph of $G$ by contracting disjoint subgraphs each with radius at most $r$. Let $G\,\triangledown\,r$ be the set of all $r$-shallow-minors of $G$, and let $\nabla_{r}(G)$$:=\max\{|E(H)|/|V(H)|\colon H\in G\,\triangledown\,r\}$. A hereditary graph class $\mathcal{G}$ has bounded expansion with expansion function $f_{\mathcal{G}}:\mathbb{N}\cup\{0\}\to\mathbb{R}$ if $\nabla_{r}(G)\leq f_{\mathcal{G}}(r)$ for every $r\geq 0$ and graph $G\in\mathcal{G}$. Bounded expansion is a robust measure of sparsity with many characterisations Zhu (2009); Nešetřil and Ossona de Mendez (2008); Nešetřil and Ossona de Mendez (2012). For example, Zhu (2009) showed that a hereditary graph class $\mathcal{G}$ has bounded expansion if there is a function $f$ such that $\operatorname{scol}_{r}(G)\leq f(r)$ for every $r\geq 1$ and graph $G\in\mathcal{G}$. Examples of graph classes with bounded expansion include classes that have bounded maximum degree Nešetřil and Ossona de Mendez (2008), bounded stack number Nešetřil et al. (2012), bounded queue-number Nešetřil et al. (2012), bounded nonrepetitive chromatic number Nešetřil et al. (2012), strongly sublinear separators Dvořák and Norin (2016), as well as proper minor-closed graph classes Nešetřil and Ossona de Mendez (2008). See the book by Nešetřil and Ossona de Mendez (2012) for further background on bounded expansion.
Given the richness of generalised colouring numbers, several attempts have been made to extend these parameters to the dense setting. In a recent breakthrough, Toruńczyk (2023) introduced cop-width and flip-width, new families of graph parameters that generalise treewidth, degeneracy, generalised colouring numbers, clique-width and twin-width. Their definitions are inspired by a game of cops and robber by Seymour and Thomas (1993):
“The robber stands on a vertex of the graph, and can at any time run at great speed to any other vertex along a path of the graph. He is not permitted to run through a cop, however. There are $k$ cops, each of whom at any time either stands on a vertex or is in a helicopter (that is, is temporarily removed from the game). The objective of the player controlling the movement of the cops is to land a cop via helicopters on the vertex occupied by the robber, and the robber’s objective is to elude capture. (The point of the helicopters is that cops are not constrained to move along paths of the graph – they move from vertex to vertex arbitrarily.) The robber can see the helicopter approaching its landing spot and may run to a new vertex before the helicopter actually lands” Seymour and Thomas (1993).
Seymour and Thomas (1993) showed that the least number of cops needed to win this game on a graph $G$ is in fact equal to $\operatorname{tw}(G)+1$, thus giving a min-max theorem for treewidth. Toruńczyk (2023) introduced the following parameterised version of this game: for some fixed $r\in\mathbb{N}$, the robber runs at speed $r$. So in each round, after the cops have taken off in their helicopters to their new positions (they may also choose to stay put), which are known to the robber, and before the helicopters have landed, the robber may traverse a path of length at most $r$ that does not run through a cop that remains on the ground. This variant is called the cop-width game with radius $r$ and width $k$, if there are $k$ cops, and the robber can run at speed $r$. For a graph $G$, the radius-$r$ cop-width of $G$, $\operatorname{copwidth}_{r}(G)$, is the least number $k\in\mathbb{N}$ such that the cops have a winning strategy for the cop-width game played on $G$ with radius $r$ and width $k$.
Say a class of graphs $\mathcal{G}$ has bounded cop-width if there is a function $f$ such that for every $r\in\mathbb{N}$ and graph $G\in\mathcal{G}$, $\operatorname{copwidth}_{r}(G)\leq f(r)$. Toruńczyk (2023) showed that bounded cop-width coincides with bounded expansion.
Theorem 1 (Toruńczyk (2023))
A class of graphs has bounded expansion if and only if it has bounded cop-width.
As such, only sparse graph classes have bounded cop-width. Flip-width is then defined as a dense analog of cop-width. Here, the cops have enhanced power where they are allowed to perform flips on subsets of the vertex set of the graph with the goal of isolating the robber. For a fixed graph $G$, applying a flip between a pair of sets of vertices $A,B\subseteq V(G)$ results in the graph obtained from $G$ by inverting the adjacency between any pair of vertices $a,b$ with $a\in A$ and $b\in B$. If $G$ is a graph and $\mathcal{P}$ is a partition of $V(G)$, then call a graph $G^{\prime}$ a $\mathcal{P}$-flip of $G$ if $G^{\prime}$ can be obtained from $G$ by performing a sequence of flips between pairs of parts $A,B\in\mathcal{P}$ (possibly with $A=B$). Finally, call $G^{\prime}$ a $k$-flip of $G$, if $G^{\prime}$ is a $\mathcal{P}$-flip of $G$, for some partition $\mathcal{P}$ of $V(G)$ with $|\mathcal{P}|\leq k$.
The flip-width game with radius $r\in\mathbb{N}$ and width $k\in\mathbb{N}$ is played on a graph $G$ by two players, flipper and runner. Initially, $G_{0}=G$ and $x_{0}$ is a vertex of $G$ chosen by the runner. In each round $i\geq 1$ of the game, the flipper announce a new $k$-flip $G_{i}$ of $G$. The runner, knowing $G_{i}$, moves to a new vertex $x_{i}$ by running along a path of length at most $r$ from $x_{i-1}$ to $x_{i}$ in the previous graph $G_{i-1}$. The game terminates when $x_{i}$ is an isolated vertex in $G_{i}$. For a fixed $r\in\mathbb{N}$, the radius-$r$ flip-width of a graph $G$, $\operatorname{flipwidth}_{r}(G)$, is the least number $k\in\mathbb{N}$ such that the flipper has a winning strategy in the flip-width game of radius $r$ and width $k$ on $G$.
In contrast to cop-width, flip-width is well-behaved on dense graphs. For example, one can easily observe that for all $r\in\mathbb{N}$, the radius-$r$ flip-width of a complete graph is equal to 1. Moreover, to demonstrate the robustness of flip-width, Toruńczyk (2023) proved the following results.
Theorem 2 (Toruńczyk (2023))
  • Every class of graphs with bounded expansion has bounded flip-width.
  • Every class of graphs with bounded twin-width has bounded flip-width.
  • If a class of graphs $\mathcal{G}$ has bounded flip-width, then any first-order interpretation of $\mathcal{G}$ also has bounded flip-width.
  • There is a slicewise polynomial algorithm that approximates the flip-width of a given graph $G$.
So flip-width is considered to be a good analog of generalised colouring numbers for dense graphs. See Toruńczyk (2023); Chang et al. (2025); Eppstein and McCarty (2023) for further results and conjectures on flip-width.
Results
In this paper, we bound the cop-width of a graph by its strong colouring numbers.
{restatable}
thmMainStrongColouring For every $r\in\mathbb{N}$, every graph $G$ has $\operatorname{copwidth}_{r}(G)\leq\operatorname{scol}_{4r}(G)$.
Previously, the best known bounds for the cop-width of a sparse graph was through its weak-colouring numbers. Toruńczyk (2023) showed that for every $r\in\mathbb{N}$, every graph $G$ has
$\operatorname{copwidth}_{r}(G)\leq\operatorname{wcol}_{2r}(G)+1.$
Moreover, if $G$ excludes $K_{t,t}$ as a subgraph, then $\operatorname{flipwidth}_{r}(G)\leq(\operatorname{copwidth}_{r}(G))^{t}.$ While graph classes with bounded strong colouring numbers have bounded weak colouring numbers, strong colouring numbers often give much better bounds than weak colouring numbers. In fact, Grohe et al. (2018) and Dvořák et al. (2022) have both shown that there are graph classes with polynomial strong colouring numbers and exponential weak colouring numbers.
{restatable}
thmLinearStrongCopFlip Every class of graphs with linear cop-width has linear flip-width. It therefore follows that that every graph class with linear strong colouring numbers has linear cop-width and linear flip-width.
Second, Section 1 gives improved bounds for the cop-width of many well-studied sparse graphs. A graph $H$ is a minor of a graph $G$ if $H$ is isomorphic to a graph that can be obtained from a subgraph of $G$ by contracting edges. A graph $G$ is $H$-minor-free if $H$ is not a minor of $G$. Van den Heuvel, Ossona de Mendez, Quiroz, Rabinovich and Siebertz van den Heuvel et al. (2017) showed that for every $r\in\mathbb{N}$, every $K_{t}$-minor-free graph $G$ has $\operatorname{scol}_{r}(G)\leq\binom{t-1}{2}(2r+1)$. So Section 1 implies the following.
{restatable}
thmKtMain For all $r,t\in\mathbb{N}$, every $K_{t}$-minor-free graph $G$ has
$\operatorname{copwidth}_{r}(G)\leq\binom{t-1}{2}(8r+1).$
By Section 1, $K_{t}$-minor-free graphs also have linear flip-width; see Corollary 5 for an explicit bound. In regard to the previous best known bound for this class of graphs, van den Heuvel et al. van den Heuvel et al. (2017) showed that for every $r\in\mathbb{N}$, every $K_{t}$-minor-free graph $G$ has $\operatorname{wcol}_{r}(G)\in O_{t}(r^{t})$. By the aforementioned result of Toruńczyk (2023), the previous best known bound for the cop-width and flip-width of $K_{t}$-minor-free graph $G$ was:
$\operatorname{copwidth}_{r}(G)\in O_{t}(r^{t-1})\text{ and }\operatorname{flipwidth}_{r}(G)\in O_{t}(r^{(t-1)^{2}}).$
Section 1 is also applicable to non-minor-closed graph classes. For a surface $\Sigma$, we say that a graph $G$ is $(\Sigma,k)$-planar if $G$ has a drawing on $\Sigma$ such that every edge of $G$ is involved in at most $k$ crossings. A graph is $(g,k)$-planar if it is $(\Sigma,k)$-planar for some surface $\Sigma$ with Euler genus at most $g$. Such graphs are widely studied (Pach and Tóth, 1997; Dujmović et al., 2023, 2017a, 2017b) and are a classic example of a sparse non-minor-closed class of graphs.
Van den Heuvel and Wood van den Heuvel and Wood (2017, 2018) showed that for every $r\in\mathbb{N}$, every $(g,k)$-planar graph $G$ has $\operatorname{scol}_{r}(G)\leq(4g+6)(k+1)(2r+1)$. So Section 1 implies the following.
Theorem 3
For all $g,k,r\in\mathbb{N}$, every $(g,k)$-planar graph $G$ has
$\operatorname{copwidth}_{r}(G)\leq(4g+6)(k+1)(8r+1).$
See Hickingbotham and Wood (2024); van den Heuvel and Wood (2017); van den Heuvel et al. (2017); Dvořák et al. (2021) for other graph classes that Section 1 applies to.
2 Proofs
We now prove our main theorem.
[…]
Proof 2.1.
Let $n:=|V(G)|-1$ and let $(v_{0},v_{1},\dots,v_{n})$ be a total order $\preceq$ of $V(G)$ where $|R(G,\preceq,v,4r)|\leq\operatorname{scol}_{4r}(G)$ for every $v\in V(G)$. For every $s\in\mathbb{N}$ and $v_{i},v_{j}\in V(G)$ where $i\leq j$, let ${M(v_{i},v_{j},s)}$ be the set of vertices $w\in V(G)$ for which there is a path $v_{j}=w_{0},w_{1},\dots,w_{s^{\prime}}=w$ of length $s^{\prime}\in[0,s]$ such that $w\preceq v_{i}$ and $v_{i}\prec w_{\ell}$ for all $\ell\in[s^{\prime}-1]$.
Claim: For all $v_{i},v_{j}\in V(G)$ where $i\leq j$, $|M(v_{i},v_{j},2r)|\leq\operatorname{scol}_{4r}(G)$.
Proof 2.2.
Let $k\in[i,j]$ be minimal such that $v_{k}\in Q(G,\preceq,v_{j},2r)$. So $G$ contains a path $P=(v_{j}=u_{0},\dots,u_{r^{\prime}}=v_{k})$ of length $r^{\prime}\in[0,2r]$ such that $v_{k}\prec u_{\ell}$ for all $\ell\in[r^{\prime}-1]$. We claim that $M(v_{i},v_{j},2r)\subseteq R(G,\preceq,v_{k},4r)$. Let $w\in M(v_{i},v_{j},2r)$. Then there is a path $P^{\prime}=(v_{j}=w_{0},\dots,w_{s^{\prime}}=w)$ of length $s^{\prime}\in[0,2r]$ such that $w\preceq v_{i}$ and $v_{i}\prec w_{\ell}$ for all $\ell\in[s^{\prime}-1]$. Suppose there is an $\ell\in[s^{\prime}-1]$ such that $w_{\ell}\prec v_{k}$. Choose $\ell$ to be minimum. Then $w_{\ell}\in Q(G,\preceq,v_{j},2r)$ since $w_{\ell}\prec v_{k}\preceq w_{a}$ for each $a\in[0,\ell-1]$, which contradicts the choice of $k$. So $v_{k}\preceq w_{\ell}$ for all $\ell\in[s^{\prime}-1]$. By taking the union of $P$ and $P^{\prime}$, it follows that $G$ contains a $(v_{k},w)$-walk $W$ of length at most $4r$ such that $v_{k}\prec u$ for all $u\in V(W)\setminus\{v_{k},w\}$. Therefore $w\in R(G,\preceq,v_{k},4r)$ and so $|M(v_{i},v_{j},2r)|\leq|R(G,\preceq,v_{k},4r)|\leq\operatorname{scol}_{4r}(G).$
For each round $i\geq 0$ until the robber is caught, we will define a tuple $(v_{i},x_{i},C_{i},D_{i},V_{i},P_{i})$ where:
  1. (1)
    $x_{i}\in V(G)$ is the location of the robber at the end of round $i$;
  2. (2)
    $C_{0}:=\{v_{0}\}$ and $C_{i}:=M(v_{i},x_{i-1},2r)$ is the set of vertices that the cops are on at the end of round $i$ for each $i\geq 1$;
  3. (3)
    $D_{0}:=\emptyset$ and $D_{i}:=C_{i-1}\cap C_{i}$ is the set of vertices where cops remain put throughout round $i$ for each $i\geq 1$;
  4. (4)
    $V_{i}:=\{v_{0},\dots,v_{i}\}$ where $v_{i}$ is defined by the total order $\preceq$ of $V(G)$; and
  5. (5)
    $P_{0}:=\emptyset$ and $P_{i}$ is the $(x_{i-1},x_{i})$-path of length at most $r$ that the robber traverses during round $i$ for each $i\geq 1$.
Note that we require $v_{i}\preceq x_{i-1}$ in order for $M(v_{i},x_{i-1},2r)$ to be well-defined. We will prove that the tuple is indeed well-defined by inductively maintain the following invariants for each round $i\geq 0$:
  1. (6)
    $v_{i}\preceq x_{i}$;
  2. (7)
    if $i\geq 1$, every path in $G$ from $x_{i-1}$ to a vertex in $V_{i-1}$ of length at most $r$ contains a vertex from $D_{i}$;
  3. (8)
    $M(v_{i},x_{i},r)\subseteq C_{i}$;
  4. (9)
    if $i\geq 1$, then $V(P_{i})\cap V_{i-1}=\emptyset$; and
  5. (10)
    if $v_{i}=x_{i}$, then the robber is caught.
Together with the previous claim, 2, 6 and 10 imply that the robber is caught within $n$ rounds using at most $\operatorname{scol}_{4r}(G)$ cops.
Initialise the game of cops and robber with the robber on some vertex $x_{0}$ in $V(G)$, one cop on $v_{0}$ and the remaining cops all in the helicopters. Define the tuple $(v_{0},x_{0},C_{0},D_{0},V_{0},P_{0})$ according to 1, 2, 3, 4 and 5. Clearly such a tuple is well-defined. Moreover, it is easy to see that the tuple satisfies 6, 7, 8, 9 and 10.
Now suppose we are at round $i\geq 1$ and the robber has not yet been caught. By induction, we may assume that there is a tuple $(v_{i-1},x_{i-1},C_{i-1},D_{i-1},V_{i-1},P_{i-1})$ for round $i-1$ which satisfies 1, 2, 3, 4, 5, 6, 7, 8, 9 and 10. Since the robber has not yet been caught, 6 and 10 imply that $v_{i-1}\prec x_{i-1}$, so $v_{i}\preceq x_{i-1}$. Therefore, there is a well-defined tuple $(v_{i},x_{i},C_{i},D_{i},V_{i},P_{i})$ which satisfies 1, 2, 3, 4 and 5. We now show that $(v_{i},x_{i},C_{i},D_{i},V_{i},P_{i})$ satisfies the additional invariants.
We first verify 7. Let $F_{i}:=M(v_{i-1},x_{i-1},r)$. Let $u\in V_{i-1}$ and suppose there is a path $P^{\star}=(x_{i-1}=w_{0},w_{1},\dots,w_{r^{\prime}}=u)$ in $G$ where $r^{\prime}\in[0,r]$. Consider the minimal $j\in[r^{\prime}]$ such that $w_{j}\in V_{i-1}$. Since $\{w_{1},\dots,w_{j-1}\}\cap V_{i-1}=\emptyset$, it follows that $w_{j}\in F_{i}$. So for every $u\in V_{i-1}$, every $(x_{i-1},u)$-path in $G$ of length at most $r$ contains a vertex from $F_{i}$. By 2 and 8 (from the $i-1$ case), it follows that $F_{i}\subseteq C_{i-1}\cap C_{i}$. So 7 follows from 3. Now since the robber is not allowed to run through a cop that stays put, 9 follows by 3, 5 and 7. Property 6 then immediately follows from 9 since $x_{i}\in V(P_{i})$. Now consider a vertex $y\in M(v_{i},x_{i},r)$. Then $G$ contains a path $P^{\prime}=(x_{i}=w_{0},w_{1},\dots,w_{r^{\prime}}=y)$ of length $r^{\prime}\in[0,r]$ such that $v_{i}\preceq w_{j}$ for all $j\in[r^{\prime}-1]$. By taking the union of $P^{\prime}$ and $P_{i}$, it follows that $G$ contains an $(x_{i-1},y)$-walk $W$ of length at most $2r$. Moreover, by 7 and the definition of $P^{\prime}$, $v_{i}\preceq z$ for all $z\in V(W)\setminus\{x_{i-1},y\}$. So $y\in M(v_{i},x_{i-1},2r)$ and thus 2 implies 8. Finally, if $v_{i}=x_{i}$ then 8 implies $x_{i}\in C_{i}$, so 10 follows by 1 and 2, as required.
To prove Section 1, we leverage known results concerning neighbourhood diversity. Neighbourhood diversity is a well-studied concept with various applications Reidl et al. (2019); Eickmeyer et al. (2017); Gajarský et al. (2017); Paszke and Pilipczuk (2020); Bonnet et al. (2022, 2024); Joret and Rambaud (2024). Let $G$ be a graph. For a set $S\subseteq V(G)$, let ${\pi_{G}(S)}:=|\{N_{G}(v)\cap S\colon v\in V(G)\setminus S\}|$. For $k\in\mathbb{N}$, let ${{\pi_{G}(k)}:=\max\{\pi_{G}(S)\colon S\subseteq V(G),|S|\leq k\}}$.
Lemma 4.
For all $k,r\in\mathbb{N}$, every graph $G$ with $\operatorname{copwidth}_{r}(G)\leq k$ has
${\operatorname{flipwidth}_{r}(G)\leq\pi_{G}(k)+k}.$
Proof 2.3.
We claim that for every set $S\subseteq V(G)$ where $|S|\leq k$, there is a $(\pi_{G}(k)+k)$-flip that isolates $S$ while leaving $G-S$ untouched. Let $\mathcal{P}$ be a partition of $V(G)$ that partition $S$ into singleton and vertices in $v\in V(G)\setminus S$ according to $N_{G}(v)\cap S$. Then $|\mathcal{P}|\leq\pi_{G}(k)+k$. Moreover, every vertex $s\in S$ can be isolated by flipping $\{s\}$ with every part of $\mathcal{P}$ that is complete to $\{s\}$. Thus, a winning strategy for the cops in the cop-width game with radius $r$ and width $k$ can be transformed into a winning strategy for the flipper in the flip-width graph with radius $r$ and width $\pi_{G}(k)+k$, as required.
Reidl et al. (2019) showed that for every graph class $\mathcal{G}$ with bounded expansion, there exists $c>0$ such that $\pi_{G}(k)\leq ck$ for every $G\in\mathcal{G}$. Since graph classes with linear cop-width have bounded expansion, Lemma 4 imply Section 1. As a concrete example, Bonnet et al. (2022) showed that for every $K_{t}$-minor-free graph $G$ and for every set $A\subseteq V(G)$,
$\pi_{G}(A)\leq 3^{2t/3+o(t)}|A|+1.$
So Sections 1 and 4 imply the following.
Corollary 5.
For all $r,t\in\mathbb{N}$, every $K_{t}$-minor-free graph $G$ has
$\operatorname{flipwidth}_{r}(G)\leq 3^{(2t/3+o(t))}t^{2}r.$
Acknowledgement
This work was initiated at the 10th Annual Workshop on Geometry and Graphs held at Bellairs Research Institute in February 2023. Thanks to Rose McCarty for introducing the author to flip-width and to David Wood for helpful conversations. The author was supported by an Australian Government Research Training Program Scholarship when he completed this work.